# I INTRODUCTION

### v1

@@ I INTRODUCTION
To confront open-world variability, a common approach is to construct open-vocabulary 3D scene graphs with MLLMs. Class-agnostic segmentation (e.g., Segment Anything, Tokenize Anything (TAP)) yields fine-grained segments where each segment is embedded with feature representations from models such as CLIP, and objects are formed by grouping segments using a combination of weighted geometric overlap and semantic similarity. However, merging 3D primitives into object-level representations remains insufficient for open-ended tasks with variable granularity. In warehouse settings, the same query (e.g., “fragile box”) may require global aggregation, such as identifying multiple pallets stacked with fragile boxes during inbound logistics, or fine-grained resolution, such as selecting a graspable box from the top of a stacked pallet during order fulfillment. Fixed object-level retrieval therefore either collapses global context or fails to support local manipulability. While Clio [8] advances task-driven scene-graph construction, it relies on a predefined task list and repeatedly re-aggregates all primitives for each list, raising concerns about computational efficiency and scalability as primitive counts grow and tasks evolve. Moreover, most existing scene-graph frameworks remain largely object-centric and static, leaving rich inter-object spatial relations, dynamic events, and variable information granularity insufficiently addressed.
@@ II Related works
Task-Driven Representations. Another promising direction is task-shaped memory. Clio [8] applies the Information Bottleneck to cluster 3D primitives based on a predefined task set, aligning map granularity to task needs; however, modifying the task list requires reclustering all primitives, limiting scalability. ASHiTA [3] employs hierarchical IB to decompose instructions into sub-tasks and associate them with relevant 3D primitives, but still assumes an explicit task structure. In contrast, our approach (i) requires no predefined task list, supporting open-ended queries, and (ii) performs task-conditioned aggregation on only a subset of primitives, substantially reducing computation while maintaining accuracy, enabling scalability.
@@ III Problem Formulation
  • Scalable task-conditioned retrieval: Given a growing memory $M_{1:T}$, how can robots extract an optimal subset $R^{*}\subseteq M_{1:T}$ that is task-relevant, information-rich, and non-redundant—yet remains information-equivalent to the full memory w.r.t. the query—while keeping retrieval efficient as $T$ increases.
@@ V-2 WH-VQA dataset
We additionally release WH-VQA, a customized warehouse benchmark built in Isaac Sim, and expand it with 100 evaluation tasks. The dataset captures a 22-minute navigation trajectory and is more challenging than NaVQA due to the presence of many visually similar objects that require task-level disambiguation. Its task composition mirrors NaVQA—spatial (40%), binary yes/no (20%), and descriptive (20%)—while adding multimodal queries (20%) to emulate human-robot interaction with underspecified language requiring grounding against the current visual scene.
@@ VI-2 WH-VQA dataset
WH-VQA is deliberately more challenging than Na: questions demand stronger contextual reasoning and adaptation to variable semantic granularity and overall SR drops for all methods. Even in this setting, our approach performs best (Table II), roughly doubling the success rate over the object-caption baseline, clearly outperforming ReMEmbR, and reducing spatial error by about 50–60%. We also achieve the highest Recall across the Binary, Text, and Multi-modal categories, indicating more accurate and comprehensive memory retrieval.

### v2

@@ I INTRODUCTION
To confront open-world variability, a common approach is to construct open-vocabulary 3D scene graphs with large foundation models. Class-agnostic segmentation[5, 4] (e.g., Segment Anything, Tokenize Anything (TAP)) yields fine-grained segments where each segment is embedded with feature representations from models such as CLIP, and objects are formed by grouping segments using a combination of weighted geometric overlap and semantic similarity. However, merging 3D primitives into object-level representations remains insufficient for open-ended tasks with variable granularity. For example, in a warehouse, a request such as “retrieve a box from shelf with a fragile label” may correspond to many visually similar boxes distributed across staging areas, shelves, and unloading zones; moreover, dozens of such boxes may be stacked on a pallet. If retrieval uses “object” as the smallest unit, the top-k results can easily collapse into a single local region (e.g., one pallet), losing broader environmental context that is critical for disambiguation. While Clio [8] represents an important step toward task-driven scene-graph construction, it relies on a predefined task list and repeatedly re-aggregates all primitives for each specification, which may introduce additional computational overhead as primitive counts grow and task requirements evolve. Moreover, most existing scene-graph frameworks remain largely object-centric and static, leaving rich inter-object spatial relations, dynamic events, and variable information granularity insufficiently addressed.
@@ II Related works
Task-Driven Representations. Another promising direction is task-shaped memory. Clio [8] applies IB to cluster 3D primitives based on a predefined task set, aligning map granularity to task needs; however, modifying the task list requires reclustering all primitives, limiting scalability. ASHiTA [3] employs hierarchical IB to decompose high-level instructions into sub-tasks and associate them with relevant 3D primitives. Our approach (i) requires no predefined task list, supporting open-ended queries, and (ii) performs task-conditioned aggregation on only a subset of primitives, substantially reducing computation while maintaining accuracy, enabling scalability.
@@ III Problem Formulation
  • Scalable task-conditioned retrieval: Given a growing memory $M_{1:T}$, how can robots extract an optimal subset $R^{*}\subseteq M_{1:T}$ that is task-relevant, information-rich, and non-redundant, yet remains information-equivalent to the full memory w.r.t. the query, while keeping retrieval efficient as $T$ increases.
@@ V-2 WH-VQA dataset
We additionally create WH-VQA, a customized warehouse benchmark built in Isaac Sim, and expand it with 100 evaluation tasks. The dataset captures a 22-minute navigation trajectory and is more challenging than NaVQA due to the presence of many visually similar objects that require task-level disambiguation. Its task composition mirrors NaVQA—spatial (40%), binary yes/no (20%), and descriptive (20%)—while adding multimodal queries (20%) to emulate human-robot interaction with underspecified language requiring grounding against the current visual scene.
@@ VI-2 WH-VQA dataset
WH-VQA is deliberately more challenging than NaVQA: questions demand stronger contextual reasoning and adaptation to variable semantic granularity and overall SR drops for all methods. Even in this setting, our approach performs best (Table II), roughly doubling the success rate over the object-caption baseline, clearly outperforming ReMEmbR, and reducing spatial error by about 50–60%. We also achieve the highest Recall across the Binary, Text, and Multi-modal categories, indicating more accurate and comprehensive memory retrieval.